A MAN IS SUFFERING FROM HAEMOPHILIA AND CYSTIC FIBROSIS. HIS FATHER IS HETEROZYGOUS FOR CYSTIC FIBROSIS. THE PROBABILITY OF A MAN`S SPERM HAVING RECESSIVE X LINKED AS WELL AS AUTOSOMAL ALLELE IS
​1. 1/4
2. 1/16
​3. 1/2
​4. 1/8

Grade:12th pass

1 Answers

Simran Vinaik
5 years ago
There is 100% chance of receiving cystic fibrosis allele as Cystic fibrosis is an autosomal recessive disease i.e both the alleles of gene should be in recessive form to cause a disease.But 50% chances of receiving the hemophilic allele. so= 1*1/2=1/2 is the probability.